# IssueGuard: Real-Time Secret Leak Prevention Tool for GitHub Issue Reports

### v1

@@ Abstract.
GitHub and GitLab are widely used collaborative platforms whose issue-tracking systems contain large volumes of unstructured text, including logs, code snippets, and configuration examples. This creates a significant risk of accidental secret exposure, such as API keys and credentials, yet these platforms provide no mechanism to warn users before submission. We present IssueGuard, a tool for real-time detection and prevention of secret leaks in issue reports. Implemented as a Chrome extension, IssueGuard analyzes text as users type and combines regex-based candidate extraction with a fine-tuned CodeBERT model for contextual classification. This approach effectively separates real secrets from false positives and achieves an F1-score of 92.70% on a benchmark dataset, outperforming traditional regex-based scanners. IssueGuard integrates directly into the web interface and continuously analyzes the issue editor, presenting clear visual warnings to help users avoid submitting sensitive data. The source code is publicly available at https://github.com/nafiurahman00/IssueGuard , and a demonstration video is available at https://youtu.be/kvbWA8rr9cU .
@@ 4. Conclusion and Future Scope
Appendix A Tool Demonstration Walkthrough (Appendix)
This appendix describes the walkthrough used for the live demonstration of IssueGuard.
A.1. System Requirements
To successfully deploy the tool, the following environment is required:
  • Python 3.12.0 or higher
  • Google Chrome browser
A.2. Installation and Setup
A.2.1. Clone the Repository
First, clone the project repository and navigate into the directory:
git clone https://github.com/nafiurahman00/IssueGuard.git
cd IssueGuard
A.2.2. Install Python Dependencies
Install the required dependencies for the FastAPI backend:
pip install -r requirements_fastapi.txt
A.2.3. Download the Model
The pre-trained CodeBERT model requires manual installation from the GitHub releases:
  1. (1)
    Navigate to the GitHub Releases page of the repository.
  2. (2)
    Download the models.zip file from the latest release.
  3. (3)
    Extract the models.zip archive into the root directory of the project.
  4. (4)
    Verify that the model files are present at the following path: models/balanced/microsoft_codebert-base_complete/
A.2.4. Start the FastAPI Server
Launch the backend server to handle inference requests:
python main.py
A.2.5. Start the Tool Using Docker (Alternative)
As an alternative to installing dependencies locally, the backend can be started using Docker Compose:
docker compose up --build
A.2.6. Install the Chrome Extension
  1. (1)
    Open Google Chrome and navigate to chrome://extensions/.
  2. (2)
    Enable Developer mode” using the toggle switch in the top right corner.
  3. (3)
    Click on Load unpacked” and select the IssueGuardExtension directory from the cloned repository.
  4. (4)
    The IssueGuard extension should now appear in the list of active extensions.
A.3. Usage
  1. (1)
    Ensure the backend server is running locally.
  2. (2)
    Open any GitHub or GitLab issue creation page (e.g., visiting https://github.com/<username>/<project>/issues/new or visiting https://gitlab.com/<username>/<project>/-/issues/ and clicking on ”New Item”).
  3. (3)
    Start writing or pasting text into the issue description box.
IssueGuard will automatically analyze the content in real-time. Detected secrets are highlighted, and a tooltip lists all true secrets identified by the model. Candidates captured by regex that are classified as false positives are automatically ignored.
A.4. Links

### v2

@@ Abstract.
GitHub and GitLab are widely used collaborative platforms whose issue-tracking systems contain large volumes of unstructured text, including logs, code snippets, and configuration examples. This creates a significant risk of accidental secret exposure, such as API keys and credentials, yet these platforms provide no mechanism to warn users before submission. We present IssueGuard, a tool for real-time detection and prevention of secret leaks in issue reports. Implemented as a Chrome extension, IssueGuard analyzes text as users type and combines regex-based candidate extraction with a fine-tuned CodeBERT model for contextual classification. This approach effectively separates real secrets from false positives and achieves an F1-score of 92.70% on a benchmark dataset, outperforming traditional regex-based scanners. IssueGuard integrates directly into the web interface and continuously analyzes the issue editor, presenting clear visual warnings to help users avoid submitting sensitive data. The source code is publicly available at https://github.com/disa-lab/IssueGuard , and a demonstration video is available at https://youtu.be/kvbWA8rr9cU .